OpenStudy (anonymous):

After 1.00 ml of water is completely vaporized to gas, how many milliliters of vapor are produced at standard temperature and pressure? (Density of water is 1.00 g/mL; molar volume of any gas at STP is 22.4 L/mol.)

OpenStudy (aaronq):

d=m/V 1 g/mL = xg/1mL x=1g PV=nRT find number of moles n=(1g)/(18g/mol)= 0.05555 g/mol T= 25 celsius = 298.15 K P= 1 atm r= 0.08206 Latm/Kmol find V